Emma is eight years old her sister chloe is twice as old as Emma when emma is twenty-five how old will chloe be?
Found 3 solutions by richwmiller, josgarithmetic, greenestamps:
Answer by richwmiller(17219)

Emma is eight years old her sister chloe is twice as old as Emma when emma is twenty-five how old will chloe be?
Emma is eight years old and her sister Chloe is twice as old as Emma. When Emma turns twenty-five, how old will Chloe be?
e=8
c=2*8=16
16-8=8 years apart
Cloe is 8 years older than Emma
25+8=?

Answer by josgarithmetic(39616)

Write the description using correct sentence-writing mechanics.

---------------------------------------------------------
Emma is eight years old.
Her sister Chloe is twice as old as Emma.
When Emma is twenty-five, how old will Chloe be?
----------------------------------------------------------

Answer by greenestamps(13198)



If Emma is 8 and Chloe is twice as old, what is the difference in their ages?

That difference in their ages is always the same; how old will Chloe be when Emma is 25?